[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Feb 17, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Feb 17 14:00:57 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:

HSU:

Other items:
1. At 0430Z this morning, CenturyLink noticed problems with NHC's secondary  
internet circuit
(ETH1000C2-17095291...linkStatus down). This was determined to be a false  
alarm.

2. GFS model science evaluations were due back to EMC by February 3. The  
EMC Director's brief is February 17 (today) and NCEP Director's brief is  
February 21 (Tuesday). The 30-day evaluation of P-Surge should be starting  
in the next 2-3 weeks.

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170217/38691293/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list